# Understanding News Thumbnail Representativeness by Counterfactual Text-Guided Contrastive Language-Image Pretraining

### v2

@@ 5.2.1 Counterfactual text generation
(1) Candidate token set construction: The first step aims to select tokens that likely refer to the subject of a news event. A part-of-speech tagger and named entity recognizer are applied to $T$, which returns a candidate token set $C$. We tested several strategies targeting named entity categories that are frequently used in news articles, such as person, organization, and GPE [^4] Countries, cities, states.
@@
(3) Masked token prediction: Using a BERT-like masked language model Devlin et al. (2019), we predict new tokens for the masked positions in $T$. To avoid reconstructing the original text, we divide the logit scores of the output vector by a temperature parameter and repeatedly sample from the softmax-normalized distribution until a token different from the original is generated. This step produces $\tilde{T}$, which represents a synthetic event wherein the subject is modified from the original news event, yet it preserves the overarching topic of $T$. Thus, $\tilde{T}$ can be deemed a hard negative sample in contrastive learning, which diverges from $I$ but remains more closely aligned with $T$ than a random text or in-batch negatives.
In Section 6.3, we present the ablation experiments to show the effectiveness of the proposed counterfactual text generation. Additionally, we demonstrate several counterfactual text examples generated by the proposed method in Appendix C.2.
@@ 5.2.2 Training objective
We train a CLIP-like bi-encoder model to minimize
@@
Minimizing Eq. 2 aligns the text representation with the image representation by increasing the similarity of positive pair $(I_{i},T_{i})$ compared to those of negative pairs $(I_{i},T_{j})$ and $(I_{i},\tilde{T}_{i})$. Since $\tilde{T}$ represents a distinct subject against $T$, the objective would help capture whether the thumbnail image represents the news subjects in $T$.
@@ 5.2.3 Model architecture
Figure 4 depicts the neural architecture of our proposed method. We initialized the model by adopting a pretrained CLIP checkpoint [^5] https://huggingface.co/openai/clip-vit-large-patch14. The image encoder $f_{image}$ is the ViT-L/14 vision transformer (24 layers), and the text encoder $f_{text}$ is a causal text transformer (12 layers). The image and text encoder are followed by an MLP pooler mapping the input to a 768-dimensional vector, respectively. $T$ and $\tilde{T}$ are fed into the shared-weight text encoder, separately. According to the hyperparameter optimization experiment, we froze 11 out of 12 text transformer layers during training. All vision layers were updated.
@@ 5.2.4 Pretraining corpus
We used two pretraining corpora. First, we used the training split of NELA unlabeled data. Inspired by a prior study that replicated the CLIP pretraining using noisy web data Schuhmann et al. (2022), we applied the data sanitization process to the training split by filtering only those with a cosine similarity of 0.28 or higher over CLIP embeddings. The resulting 105,737 pairs for training have a high likelihood of a representative thumbnail ($L=1$). Since the body text in the NELA corpus is incomplete, we relied on the news title as the reference text for pretraining. Second, inspired by the recent research on the importance of high-quality samples for pretraining Zhou et al. (2023); Schuhmann et al. (2022), we used a high-quality news corpus published by the BBC, a UK-based news source rated as trustworthy by MBFC Verma et al. (2023). This corpus comprises 196,538 pairs, each containing a news title, an editor-written summary, and a thumbnail image. The corpus encompasses news articles released through the official website from 2009 to 2021. We applied the sanitation process to the dataset in the same manner as with the NELA corpus, leaving 58,210 samples for contrastive training. We ensured no overlap between the BBC training set and the NELA validation/test splits. Since NewsTT originates from the NELA corpus, we used the NELA validation split for hyperparameter optimization. In the experiments using the BBC corpus, we tested two variants using the news title and human-written summary as reference text.
Figure 4: Neural architecture of CFT-CLIP
6 Evaluation Experiments
@@ 6.1 Experimental setups
We employed the f1 score and the Spearman rank correlation coefficient to evaluate binary prediction and cosine similarity scores, respectively. All experiments were conducted using five different random seeds, and we reported the average performance along with the standard error. The t-test was used for estimating statistical significance. We used the SpaCy pipeline for named entity recognition. A BERT-base checkpoint was used for the masked language prediction. All checkpoints used for experiments and implementation details can be found in Appendix A. We optimized all hyperparameters using a validation split that included 10,000 examples from NELA.
As baselines, we measured the ability of pretrained vision and language models for comparison with the proposed methods for the target task. In addition to CLIP, we used BLIP and BLIP-2, which are a family of vision language encoder that is based on bootstrapped training Li et al. (2022b, 2023). They achieved the state-of-the-art performance for the vision-language benchmarks. BLIP-2$+$SBERT is a pipelined approach that integrates BLIP-2 with Sentence BERT. By obtaining a caption text from an input image $I$ by BLIP-2, we assessed its semantic similarity to the reference text using SentenceBERT. To investigate the effects of domain-adaptive pretraining, we also included CLIPAdapt as a baseline model, which was trained by using the CLIP objective on the training corpus. For inference, we used the summary text as the reference text $T$ according to the results of an ablation experiment (Table A1).
@@
CFT-CLIP 0.815$\pm$0.003 0.488$\pm$0.004
CLIPAdapt 0.766$\pm$0.005 0.455$\pm$0.004
CLIP 0.763 0.408
BLIP 0.737 0.406
BLIP-2 0.708 0.418
BLIP-2 $+$ SBERT 0.694 0.339
@@
6.2 Results
Table 1 presents the model comparison results of the baseline and proposed methods. Here, we reported the performance of CFT-CLIP and CLIPAdapt models that were pretrained using the BBC dataset with the summary text. For CFT-CLIP, we used the model targeting person-labeled entity tokens. The decisions were based on the ablation experiments on the effect of pretraining corpus (Table 2) and counterfactual text generation (Table 3), respectively. Here, we make three observations. First, among the pre-trained models (the bottom 4 rows), CLIP achieved the best f1 of 0.763, and BLIP-2 was the best in terms of the Spearman coefficient with 0.418. The effectiveness of CLIP suggests that its bi-encoder architecture is suitable for tackling the task. Second, CLIPAdapt outperformed all the pretrained checkpoints. This suggests that domain-adapted continued pretraining can improve the performance for the target task, congruent with the finding of a previous study Gururangan et al. (2020). Third, CFT-CLIP outperformed all baseline methods with the f1 of 0.815 and the Spearman coefficient of 0.488 (p$<$0.001). The performance gap with CLIPAdapt, the second-best method, is significant: 0.049 for f1 and 0.033 for the Spearman coefficient.
@@ 6.3 Ablation experiments
Pretraining corpus
Table 2 presents the performance of CFT-CLIP models with varying pretraining corpus. The BBC dataset resulted in superior model performance compared to the NELA dataset, following the same distribution of the labeled dataset. Given that the BBC is recognized for upholding high journalistic standards, this observation suggests that leveraging a high-quality singular source may be more beneficial than employing articles published by diverse news sources for contrastive pretraining. This finding aligns with the recent research on large language models Zhou et al. (2023), emphasizing that the data quality is more important than the size of the training corpus. In the experiments comparing title and editor-written summaries using BBC, we could not find a clear winner. While using the summary text led to the best f1 of 0.815, using the title achieved the best Spearman coefficient of 0.501. This suggests that both news headlines and summary text can serve as a useful proxy for news content, in line with established journalism principles The Associate Press (2022). Since f1 is a more proper metric for evaluating classification ability, we used BBC with the summary text for the other experiments.
@@
BBC ($T$: summary) 0.815$\pm$0.003 0.488$\pm$0.004
BBC ($T$: title) 0.790$\pm$0.007 0.501$\pm$0.001
NELA ($T$: title) 0.771$\pm$0.003 0.444$\pm$0.002
Table 2: Effects of pretraining corpus
Target token F1 Spearman
Person 0.815$\pm$0.003 0.488$\pm$0.004
Organization 0.783$\pm$0.002 0.439$\pm$0.003
GPE 0.761$\pm$0.004 0.404$\pm$0.004
All 0.785$\pm$0.002 0.454$\pm$0.003
Random (15%) 0.715$\pm$0.013 0.461$\pm$0.005
Random (30%) 0.680$\pm$0.007 0.458$\pm$0.002
Counterfactual text generation
We evaluated the performance of CFT-CLIP variants targeting different types of entity tokens in counterfactual text generation. We specifically targeted the frequently used named entity tokens in the news text: person, organization, and GPE. We also tested the model targeting all three token types, denoted All in the table. The top 4 rows in Table 3 present the comparison results. We found that targeting person-labeled entity tokens achieved the best performance. This finding could be explained by the prevalent coverage of person entities, such as politicians, in news events Park et al. (2021b); Müller-Budack et al. (2020). Additionally, we tested baseline strategies that replace tokens at random positions for being used for negative samples in contrastive learning, as investigated in previous studies Nishikawa et al. (2022); Robinson et al. (2021). We evaluated two versions that select 15% and 30% tokens, which were predicted by the same BERT backbone as the proposed method. The bottom two rows in Table 3 show the model performance, achieving significantly worse f1 and Spearman coefficients than the CFT-CLIP variants (p$<$0.001). This finding supports the effectiveness and necessity of counterfactual text generation in the proposed method.
LLMs for counterfactual text generation
CFT-CLIP used a BERT-like masked language model to generate counterfactual text for contrastive updates. An autoregressive large language model, such as GPT Radford et al. (2018), could serve as the purpose. To test the effectiveness, we ran the OpenAI API for generating counterfactual text for the BBC news summary by GPT 3.5-Turbo. The generated counterfactual text was used for contrastive update, following the objective in Eq. 2. The used prompt is available in Appendix D.
As shown in Table 4, the GPT-based contrastive update was not as successful as the proposed method achieved by a BERT-like masked language model. The select-and-replace approach of the proposed model could generate a more suitable sample to be used for hard negatives in contrastive updates.
Model F1 Spearman
CFT-CLIP 0.815$\pm$0.003 0.484$\pm$0.004
GPT-based 0.640$\pm$0.019 0.445$\pm$0.003
@@ 6.4 Error analysis
To identify the remaining challenges, we analyzed common error categories for the sampled 100 error cases. We observed four recurring categories: entity recognition failures (46%), external knowledge required (14%), deep textual understanding required (14%), and deep visual understanding required (14%). Figure 5 presents two error examples. In the first example belonging to the entity recognition failure case, the model failed to identify Brian Sicknick, the officer involved in the George Floyd case. In the second example, the image shows Anne-Marie Trevelyan, one of the ministers of the UK government. The model could not find the link without referring to the external knowledge.
Figure 5: Error examples
7 Discussion
@@
This study has several future directions. First, according to the rule of 5Ws, this study focused on the Who aspect of news thumbnails. Future research could extend its focus to cover other aspects of 5Ws, such as whether the image represents the action of the subject, i.e., What. Question-answering approaches could be investigated, as done in a recent study Rani et al. (2023). Second, the proposed method could be improved to address the challenges identified in the error analysis (Section 6.4). Future studies could investigate the use of prompt-based visual language models, such as Flamingo Alayrac et al. (2022) or GPT-4V Achiam et al. (2023), to allow for handling the errors involving deep visual understanding. The proposed CFT-CLIP could assist with aligning the vision and language representation. Third, more broadly, this study could be extended to the multimodal misinformation research by adopting the CFT-CLIP for evidence retrieval or training multimodal classifiers Luo et al. (2021); Mishra et al. (2022); Yao et al. (2023).
8 Conclusion
Understanding the representativeness of a news thumbnail is not only for holding journalistic standards but also for building a trustworthy online environment. To tackle the critical task, this study proposes NewsTT, a paired dataset of news thumbnails and text, and CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining method. The proposed method outperformed baseline models, supporting the research hypothesis that contrasting news text with its counterfactual will enhance the cross-modal understanding ability.
@@ 8 Conclusion
This study bears several limitations. First, the scale of NewsTT is limited. Since the annotation task is complicated, we chose to do an in-lab annotation to provide high-quality labels for evaluation. Since the dataset reflects the distribution of real-world news articles and the label quality is high, it can be used for a proper evaluation corpus of the zero-shot methods. Future studies could scale up the dataset via crowdsourcing with the annotation scheme developed in this study. Second, since we continued to pretrain the CLIP encoders, CFT-CLIP inherits the weakness of the CLIP vision encoder. CLIP might be culturally biased toward the Western countries where the pretraining dataset may originate. Also, the pretrained CLIP checkpoint used in this study cannot handle the entire body text as input because the maximum token length is 77, which is smaller than the average body text length. While the use of summary text mitigates the limitation, future studies could use a text encoder that can handle a long sequence to exploit the news article fully. Third, this study focused on developing a zero-shot classifier based on a CLIP-like dual encoder, which does not involve labeled data for training. The performance could be boosted by developing a fine-tuned classifier or few-shot prompt learning methods in future studies.
@@
This study introduced CFT-CLIP, a contrastive learning framework for training the image-text multimodal encoder. For pretraining, this study used publicly available news articles shared by news media. While we tried to have a high-quality corpus for pretraining, it is possible that the model learned hidden biases in online news. Also, Since CFT-CLIP was updated from the pretrained CLIP weights, it may inherit the bias of CLIP. A user should be cautious about applying the method to problems in a general context and be aware of a potential bias. A user should acknowledge the potential harm that can be induced by applying the pretrained embedding. We have fewer privacy concerns because our study used openly accessible news data that may follow strict internal guidelines. The NELA-GT-2021 was shared under the license of CC BY-NC 4.0, and the BBC corpus was shared under the MIT license. We will share NewsTT with CC BY NC 4.0. Some of the text was edited using AI assistants, such as ChatGPT and Grammarly.
Appendix A Configuration details

### v3

@@ 5.2.1 Counterfactual text generation
(1) Candidate token set construction: The first step aims to select tokens that likely refer to the actor of a news event. A part-of-speech tagger and named entity recognizer are applied to $T$, which returns a candidate token set $C$. We tested several strategies targeting named entity categories that are frequently used in news articles, such as person, organization, and GPE [^4] Countries, cities, states.
@@
(3) Masked token prediction: Using a masked language model, e.g., BERT Devlin et al. (2019), we predict new tokens for the masked positions in $T$. To avoid reconstructing the original text, we divide the logit scores of the output vector by a temperature parameter and repeatedly sample from the softmax-normalized distribution until a token different from the original is generated. This step produces $\tilde{T}$, which represents a synthetic event wherein the subject is modified from the original news event, yet it preserves the overarching topic of $T$. Thus, $\tilde{T}$ can be deemed a hard negative sample in contrastive learning, which diverges from $I$ but remains more closely aligned with $T$ than a random text or in-batch negatives.
In Section 6.3, we present the ablation experiments to show the effectiveness of the proposed counterfactual text generation. Additionally, we demonstrate several counterfactual text examples generated by the proposed method in Appendix D.2.
@@ 5.2.2 Training objective
We train an image and text bi-encoder model to minimize
@@
Minimizing Eq. 2 aligns the text representation with the image representation by increasing the similarity of positive pair $(I_{i},T_{i})$ compared to those of negative pairs $(I_{i},T_{j})$ and $(I_{i},\tilde{T}_{i})$. Since $\tilde{T}$ represents a distinct news event of which actors were replaced from $T$, the objective would help a bi-encoder model learn to capture whether the thumbnail image represents the news actors that can be identified from $T$.
@@ 5.2.3 Model architecture
Figure 4 depicts the neural architecture used for the proposed method. We initialized the model by adopting a pretrained CLIP checkpoint [^5] https://huggingface.co/openai/clip-vit-large-patch14. The image encoder $f_{image}$ is the ViT-L/14 vision transformer (24 layers), and the text encoder $f_{text}$ is a causal text transformer (12 layers). The image and text encoder are followed by an MLP pooler mapping the input to a 768-dimensional vector, respectively. $T$ and $\tilde{T}$ are fed into the shared-weight text encoder, separately. According to the hyperparameter optimization experiment, we froze 11 out of 12 text transformer layers during training. All vision layers were updated.
Figure 4: Neural architecture of CFT-CLIP
@@ 5.2.4 Pretraining corpus
We used two pretraining corpora. First, we used the training split of NELA unlabeled data. Inspired by a prior study that replicated the CLIP pretraining based on a web data collection Schuhmann et al. (2022), we applied the data sanitization process to the training split of unlabeled data by filtering only those with a cosine similarity of 0.28 or higher over CLIP embeddings. The resulting 105,737 pairs for the unlabeled training have a high likelihood of a representative thumbnail ($L=1$). Since the body text in the NELA corpus is incomplete, we relied on the news title as the reference text for pretraining. Second, inspired by the recent research on the importance of high-quality samples for pretraining Zhou et al. (2024); Schuhmann et al. (2022), we used a high-quality news corpus published by the BBC, a UK-based news source rated as trustworthy by MBFC Verma et al. (2023). This corpus comprises 196,538 pairs, each containing a news title, an editor-written summary, and a thumbnail image. The corpus encompasses news articles released through the official website from 2009 to 2021. We applied the sanitation process to the dataset in the same manner as with the NELA corpus, leaving 58,210 samples for contrastive training. We ensured no overlap between the BBC training set and the NELA validation/test splits. Since NewsTT originates from the NELA corpus, we used the NELA validation split for hyperparameter optimization. In the experiments using the BBC corpus, we tested two variants using the news title and human-written summary as reference text.
6 Evaluation
We conducted evaluation experiments to understand the effects of the proposed method and test the research hypothesis. In Section 6.2, we examined how the pretrained and proposed vision language models perform for the target task. In ablation experiments (Section 6.3), we analyzed the effects of each module of the proposed method in more detail. Lastly, in Section 6.4, we conducted qualitative analyses to identify the error patterns where the proposed method fails.
@@ 6.1 Experimental setups
As baselines, we measured the ability of pretrained vision language models. In addition to CLIP, we used BLIP and BLIP-2, which are a family of vision language encoders that is based on bootstrapped training Li et al. (2022a, 2023). They achieved state-of-the-art performance in the vision-language benchmarks. BLIP-2$+$SBERT is a pipelined approach that integrates BLIP-2 with SentenceBERT. By obtaining a caption text from an input image $I$ by BLIP-2, we assessed its semantic similarity to the reference text using SentenceBERT. To investigate the effects of domain-adaptive pretraining, we also included CLIPAdapt as a baseline model, which was trained by using the CLIP objective on the training corpus. For inference, we used the summary text as the reference text $T$ according to the results of an ablation experiment (Table A2).
To evaluate the baseline and proposed methods, we employed the f1 score and the Spearman rank correlation coefficient to evaluate binary prediction and cosine similarity scores, respectively. All experiments were conducted using five different random seeds, and we reported the average performance along with the standard error. The t-test was used for estimating statistical significance. We used the SpaCy pipeline for named entity recognition. A BERT-base checkpoint was used for the masked language prediction. All checkpoints used for experiments and implementation details can be found in Appendix B. We optimized all hyperparameters using the NELA validation split.
@@
CFT-CLIP 0.815$\pm$0.003 0.491$\pm$0.005
CLIPAdapt 0.767$\pm$0.006 0.459$\pm$0.004
CLIP 0.763 0.409
BLIP 0.737 0.408
BLIP-2 0.707 0.415
BLIP-2 $+$ SBERT 0.694 0.341
@@
6.2 Main results
Table 2 presents the evaluation results of the baseline and proposed methods. Here, we reported the performance of CFT-CLIP and CLIPAdapt models that were pretrained using the BBC dataset with the summary text. For CFT-CLIP, we used the model targeting person-labeled entity tokens. The decisions were based on the ablation experiments on the effect of pretraining corpus (Table 5) and counterfactual text generation (Table 3), respectively. We made three observations. First, among the pre-trained models (the bottom four rows), CLIP achieved the best f1 of 0.763, and BLIP-2 was the best by the Spearman coefficient with 0.415. Given that BLIP-2 (1.17B) is 2.74 times larger than CLIP (427M), this observation suggests the effectiveness of CLIP’s bi-encoder architecture for the target task. Second, CLIPAdapt outperformed all the pretrained models. This suggests that domain-adapted continued pretraining can improve the performance for the target task, congruent with the finding of a previous study Gururangan et al. (2020). Third, CFT-CLIP outperformed all baseline methods with the f1 of 0.815 and the Spearman coefficient of 0.491 (p$<$0.001). The performance gap with CLIPAdapt, the second-best method, is significant: 0.048 for f1 and 0.032 for the Spearman coefficient. The finding suggests that incorporating the counterfactual text in its contrastive objective as hard negatives can improve the cross-modal matching ability of the vision language bi-encoder, supporting the research hypothesis.
@@ 6.3 Ablation experiments
Target token F1 Spearman
Person 0.815$\pm$0.003 0.491$\pm$0.005
Organization 0.784$\pm$0.002 0.443$\pm$0.002
GPE 0.762$\pm$0.004 0.410$\pm$0.005
All 0.785$\pm$0.003 0.457$\pm$0.003
Random (15%) 0.715$\pm$0.013 0.463$\pm$0.005
Random (30%) 0.68$\pm$0.007 0.461$\pm$0.002
What tokens should be targeted?
We evaluated the performance of CFT-CLIP variants aiming to replace different types of entity tokens in counterfactual text generation. In particular, we targeted the frequently used named entity tokens in the news text: person, organization, and GPE. We also tested the model that replaces all three token types, denoted All in the table. The top 4 rows in Table 3 present the comparison results. We found that targeting person-labeled entity tokens achieved the best performance. This finding could be explained by the prevalent coverage of person entities, such as politicians, in news events Park et al. (2021b); Müller-Budack et al. (2020). Additionally, we tested baseline strategies that replace tokens at random positions for being used for negative samples in contrastive learning, as investigated in previous studies Nishikawa et al. (2022); Robinson et al. (2021). We evaluated two variants that select 15% and 30% tokens, respectively, which were predicted by the same BERT backbone as the proposed method. The bottom two rows in Table 3 show the model performance, achieving significantly worse f1 and Spearman coefficients than the CFT-CLIP variants (p$<$0.001). According to the results, we targeted person-labeled tokens in the proposed method.
Is the masked LM necessary?
CFT-CLIP used a masked language model to generate counterfactual text for contrastive updates. An autoregressive large language model, such as GPT Radford et al. (2018), could be used alternatively. To validate the idea, we ran the OpenAI API for generating counterfactual text for the BBC news summary by GPT 3.5-Turbo. The generated text was used for contrastive update, following the objective in Eq. 2. The used prompt is available in Appendix E. As shown in Table 4, the GPT-based contrastive update was not as successful as the proposed method based on a masked language model. The proposed select-and-replace approach could generate a more suitable counterfactual sample to be used for hard negatives, rather than rewriting the whole sentence conditioned on the original text.
We also tested a baseline approach that ablates a generation model for predicting masked tokens by randomly sampling person-labeled entities from the tokens from the training set. The simple method achieved an f1 of 0.77 and a Spearman coefficient of 0.455, respectively. This finding supports the effectiveness and necessity of counterfactual text generation in the proposed method.
Model F1 Spearman
CFT-CLIP 0.815$\pm$0.003 0.491$\pm$0.005
GPT-based 0.640$\pm$0.018 0.445$\pm$0.002
@@
BBC ($T$: summary) 0.815$\pm$0.003 0.491$\pm$0.005
BBC ($T$: title) 0.790$\pm$0.007 0.504$\pm$0.001
NELA ($T$: title) 0.772$\pm$0.003 0.448$\pm$0.002
Table 5: Varying performance by the pretraining corpus
Data quality vs. quantity
Table 5 presents the performance of CFT-CLIP models with varying pretraining corpus. The BBC dataset resulted in superior model performance compared to the NELA dataset, following the same distribution of the labeled dataset. Given that the BBC is recognized for upholding high journalistic standards, this observation may indicate that leveraging a high-quality singular source is more beneficial than employing articles published by diverse news sources for contrastive pretraining. This finding is aligned with the recent research on large language models Zhou et al. (2024), emphasizing that ensuring data quality is more important than merely increasing the size of the training corpus. In the experiments comparing title and editor-written summaries using BBC, we could not find a clear winner. While using the summary text led to the best f1 of 0.815, using the title achieved the best Spearman coefficient of 0.504. This suggests that both news headlines and summary text can serve as a useful proxy for news content, in line with established journalism principles The Associate Press (2022). Since f1 is a more proper metric for evaluating classification ability, we used BBC with the summary text for the other experiments.
Figure 5: Error examples
@@ 6.4 Error analysis
To identify the remaining challenges, we analyzed error categories for the sampled 100 error cases. The first author identified the initial category by thematic coding, which were improved by the iteration of discussion with the other authors and re-annotation. We observed four recurring categories: entity recognition failures (46%), external knowledge required (14%), deep textual understanding required (14%), and deep visual understanding required (14%). Figure 5 presents two error examples. In the first example, which belongs to the entity recognition failure case, the model made a wrong prediction possibly due to the failed identification of Brian Sicknick, the officer involved in the George Floyd case. In the second example, while the image shows Anne-Marie Trevelyan, a UK minister, the model could not find the cross-modal link without referring to external knowledge.
7 Discussion and Conclusion
Automatically evaluating a news thumbnail image’s representativeness is important for holding journalistic standards and building a trustworthy online environment. To address the important but underexplored problem in the research community, this study introduced NewsTT, a paired dataset of news thumbnail images and text with high-quality labels on whether the image represents the actor of the news event. We investigated the use of vision language models for the zero-shot assessment. Our proposed CFT-CLIP outperformed larger pretrained vision language models and domain-adapted methods. This supports the research hypothesis that counterfactual news text, of which named entities are replaced by a masked language model, could enhance the cross-modal matching ability by contrastive learning.
In ablation experiments, we found that using the counterfactual text generated by an autoregressive transformer language model, GPT-3.5 Turbo, as negative samples could not achieve a better outcome than CFT-CLIP. While this supports the effectiveness of the proposed method, we do not conclude that masked language models are clear winners for the counterfactual news generation over autoregressive language models. After manually examining the results, we observed that GPT could produce plausible counterfactual news. By contrast, the generation by a masked language model sometimes led to imperfect generations, including broken grammar (Table A4). As shown in a recent study on text embedding Lee et al. (2024), an additional step might be required to ensure the quality of the generated text.
@@ 7 Discussion and Conclusion
This study has several future directions. First, according to the rule of 5Ws, this study focused on the Who aspect of news thumbnails. Future research could extend its focus to cover other aspects of 5Ws, such as whether the image represents the subject’s action, i.e., What. Question-answering approaches could be investigated, as done in a recent study Rani et al. (2023). Second, the proposed method could be improved to address the challenges identified in the error analysis (Section 6.4). Future studies could investigate the use of multimodal language models, such as InstructBLIP Dai et al. (2023), LLaVA Liu et al. (2024), or GPT-4V OpenAI (2023), to allow for handling the errors involving deep visual understanding. The proposed CFT-CLIP could assist them with aligning the vision and language representation by adopting the counterfactual text. Third, this study could be extended to the broader research on computational social science. As similarly done in previous studies Lommatzsch et al. (2022); Oostdijk et al. (2020), one could extend the target problem by addressing it as a ranking task rather than binary classification. It could enable a news application that recommends a suitable thumbnail image for a given news article. Additionally, the proposed CFT-CLIP could be adopted for automated fact verification involving image evidence Luo et al. (2021); Mishra et al. (2022); Yao et al. (2023).
@@
This study bears several limitations. First, the scale of NewsTT is limited. Since the annotation task is complicated, we chose to do an in-lab annotation to provide high-quality labels for evaluation. Since the dataset reflects the distribution of real-world news articles and the label quality is high, it can be used for a reliable evaluation corpus. Future studies could scale up the dataset via crowdsourcing with the annotation scheme developed in this study. Second, since we continued to pretrain the CLIP encoders, CFT-CLIP inherits the weakness of the CLIP vision encoder. CLIP might be culturally biased toward the Western countries where the pretraining dataset may originate. The pretrained CLIP checkpoint used in this study cannot handle the entire body text as input because the maximum token length is 77, which is smaller than the average body text length. While the use of summary text mitigates the limitation, future studies could use a text encoder that can handle a long sequence to exploit the entire news article. Third, this study focused on developing a zero-shot classifier based on a CLIP-like dual encoder, which does not involve labeled data for training. The performance could be boosted by developing a fine-tuned classifier or few-shot prompt learning methods.
@@
This study introduced CFT-CLIP, a contrastive learning framework for training an image-text multimodal encoder. For pretraining, this study used publicly available news articles shared by news media. While we tried to have a high-quality corpus for pretraining, it is possible that the model learned hidden biases in online news. Also, Since CFT-CLIP was updated from the pretrained CLIP weights, it may inherit the bias of CLIP. A user should be cautious about applying the method to problems in a general context and be aware of a potential bias. We have fewer privacy concerns because our study used openly accessible news data that may follow strict internal guidelines according to journalism principles. The NELA-GT-2021 was shared under the license of CC BY-NC 4.0, and the BBC corpus was shared under the MIT license. We will share NewsTT with CC BY NC 4.0. Some of the text was edited using AI assistants, such as ChatGPT and Grammarly.
Acknowledgements
Kunwoo Park is the corresponding author. This work was supported by Innovative Human Resource Development for Local Intellectualization program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (IITP-2024-RS-2022-00156360).
Appendix
Appendix A Unlabeled data characteristics
Table A1 shows the data characteristics for the unlabeled train dataset. In reference to Table 1, we found that the labeled data follows a similar distribution to the unlabeled NELA across various dimensions. The labeled summary text tends to have different characteristics from the unlabeled BBC summary dataset because the former was generated by GPT yet the latter is written by the news editor.
NELA BBC
Title Title Summary
Words 14.3 9.3 25.4
Nouns 3.24 2.56 5.22
Verbs 1.67 1.07 2.44
Adjectives 0.94 0.61 1.54
Named entities 2.14 1.73 2.71
Table A1: Data distribution
Appendix B Configuration details